### v1

@@ Abstract
Keywords:
Neural networks Side-channel analysis attacks OpenVINO.
@@ 1 Introduction
Disclosure. OpenAI’s ChatGPT-4 was used to improve the clarity and readability of this manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the manuscript’s content.
@@ 0 $0$8 1000
$(-1)^{b_{31}}\times 2^{b_{30}b_{29}\dot{b}_{23}-127}\times 1.b_{22}b_{21}\dots b_{0},$
@@
$1+\frac{1}{2^{b_{22}}}+\frac{1}{2^{2\times b_{21}}}\dots+\frac{1}{2^{23\times b_{0}}}.$
@@ 3.3 Correlation Power Analysis
Let $l_{i}$ denote the leakage corresponding to the input value $x_{i}$ for $i=1,2,\dots,M$. In this case, $l_{i}$ simulates the leakage of
$\boldsymbol{v}_{i}:=x_{i}\times w.$
@@
$w_{\text{guessed}}=w_{i_{0}},\quad\text{where}\quad i_{0}=\operatorname*{arg\,min}_{i:w_{i}\text{ achieves the highest absolute value of }r_{i}}\left\lvert\sum_{j=1}^{M}(\mathcal{H}_{ij}-l_{j})\right\rvert.$
@@ 4.1 Evaluation Scenarios
  • Scenario 3: The bit coefficients for the stochastic leakage model are randomly generated from a normal distribution with mean $1$, with a similar higher bit variance of $1$. In this scenario, the attacker is assumed to have accurately profiled the leakage and identified the correct bit coefficient values and, as such, can build a more precise leakage model for the key recovery attack.
@@
Using this setting, we generate a set of simulated traces. We simulate weight and bias recovery using CPA with $100,000$ traces for one simulated attack, i.e. $M=100,000$ following the notations from Section 3. For the recovery of weights, a total of $250$ attacks were conducted, with randomly generated weight values for each attack. Similarly, $250$ attacks were performed for the recovery of biases, with randomly generated bias values for each attack.
@@ 4.2 Application to Full Networks
As can be seen from Table 2, we observe that the recovered model from Scenario 1 and 3 performs almost similarly to the original GoogleNet model. This makes sense since the leakage model quite matches the leakage behavior. However, for Scenario 2, when the attacker is using the Hamming weight leakage model when targeting leakage with high variance for their bit coefficients, the performance degrades significantly.
@@ Acknowledgement
This project has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under the Programme SASPRO 2 COFUND Marie Sklodowska-Curie grant agreement No. 945478. This research is funded by the European Commission, under the Horizon Europe project aerOS, grant number 101069732.
This research is supported by the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity Research & Development Programme (Development of Secured Components & Systems in Emerging Technologies through Hardware & Software Evaluation < NRF-NCR25-DeSNTU-0001 >). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the view of National Research Foundation, Singapore and Cyber Security Agency of Singapore.

### v2

@@ 0 $0$8 1000
$(-1)^{b_{31}}\times 2^{b_{30}b_{29}\dots b_{23}-127}\times 1.b_{22}b_{21}\dots b_{0},$
@@
$1+\frac{b_{22}}{2}+\frac{b_{21}}{2^{2}}\dots+\frac{b_{0}}{2^{23}}.$
@@ 3.3 Correlation Power Analysis
Let $l_{j}$ denote the leakage corresponding to the input value $x_{j}$ for $j=1,2,\dots,M$. In this case, $l_{j}$ simulates the leakage of
$\boldsymbol{v}_{j}:=x_{j}\times w.$
@@
$w_{\text{guessed}}=w_{i_{0}},$
where
$i_{0}=\operatorname*{arg\,min}_{i:w_{i}\text{ achieves the highest absolute value of }r_{i}}\left\lvert\sum_{j=1}^{M}(\mathcal{H}_{ij}-l_{j})\right\rvert.$
An algorithmic representation of the described procedure using Hamming weight leakage model for the attack is provided in Algorithm 1.
@@
Input: $\alpha_{s}$ $(s=0,1,\dots,32)$, $w$, $w_{1},w_{2},\dots,w_{N}$, $M$, $\mu$, $\sigma^{2}$ // $\alpha_{s}$ represents the coefficients of the stochastic leakage model, $w$ denotes the correct weight value, and $M$ is the number of random inputs to be generated. $w_{1},w_{2},\dots,w_{N}$ are all the possible values of the weight. The parameters $\mu$ and $\sigma^{2}$ correspond to the mean and variance, respectively, of the normal distribution used to model noise.
1
Output: recovered weight
2 randomly generate $M$ inputs $x_{1},x_{2},\dots,x_{M}$;
array of size $M\quad$ $l$// array to store simulated leakages
3 for $j=1,2,\dots,M$ do
       $\boldsymbol{v}=x_{j}\times w$// compute the correct product
       generate noise $\varepsilon$// generate noise from a normal distribution with mean $\mu$ and variance $\sigma^{2}$
       $l[j-1]=\mathcal{L}(\boldsymbol{v})$// simulate leakage using the coefficients $\alpha_{s}$ and noise $\varepsilon$ following Equation 1
4
$\overline{l}=\displaystyle\frac{1}{M}\sum_{j=0}^{M-1}l[j]$// the averaged leakage
array of size $N\quad$ $r_{abs}$// array to store absolute values of correlation coefficients
array of size $N\quad$ $\texttt{dif}_{abs}$// array to store absolute values of differences between hypothetical leakages and the simulated leakages
5 for $i=1,2,\dots,N$ do
       array of size $M\quad$ $\hat{\boldsymbol{v}}$// array to store hypothetical product values
[…]
6       for $j=1,2,\dots,M$ do
             $\hat{\boldsymbol{v}}[j-1]=w_{i}\times x_{j}$// the hypothetical product for hypothesis of the weight value $w_{i}$ and input $x_{j}$
             $\mathcal{H}[j-1]=\text{HW}(\hat{\boldsymbol{v}}[j-1])$// the hypothetical leakage for hypothesis of the weight value $w_{i}$ and input $x_{j}$
7
      ¯$\overline{\mathcal{H}}=\displaystyle\frac{1}{M}\sum_{j=0}^{M-1}\mathcal{H}[j]$// averaged hypothetical leakage
       $r_{abs}[i-1]=\displaystyle\frac{\sum_{j=0}^{M-1}(\mathcal{H}[j]-\overline{\mathcal{H}})(l[j]-\overline{l})}{\sqrt{\sum_{j=0}^{M-1}(\mathcal{H}[j]-\overline{\mathcal{H}})^{2}}\sqrt{\sum_{j=0}^{M-1}(l[j]-\overline{l})^{2}}}$// correlation coefficient for weight hypothesis $w_{i}$
       $r_{abs}[i-1]=\lvert r_{abs}[i-1]\rvert$// compute the absolute value of the correlation coefficient
       $\texttt{dif}_{abs}[i-1]=\left\lvert\displaystyle\sum_{j=0^{M-1}}(\mathcal{H}[j]-l[j])\right\rvert$// the absolute value of the sum of the differences between the hypothetical and simulated leakage values
8
Algorithm 1 A simulated attack for recovering a single weight using the Hamming weight model
[…]
ind=indices of max_cor in $\texttt{ind}=\text{indices of {max\_cor} in }r_{abs}$// find all weights that achieve the highest absolute correlation coefficient
20 if len(ind)==1==1= = 1 then
       return $w_{\texttt{ind}[0]}$// if there is just one weight that achieve the highest absolute value of correlation coefficient, we take it to be the guessed weight
21
22else
[…]
23       return $w_{\texttt{ind2}[0]}$
@@ 4.1 Evaluation Scenarios
  • Scenario 3: The bit coefficients for the stochastic leakage model are randomly generated from a normal distribution with mean $1$, with a similar higher variance of $1$. In this scenario, the attacker is assumed to have accurately profiled the leakage and identified the correct bit coefficient values and, as such, can build a more precise leakage model for the key recovery attack.
@@
Using this setting, we generate a set of simulated traces. We simulate weight and bias recovery using CPA with $100,000$ traces for one simulated attack, i.e. $M=100,000$ following the notations from Section 3. Specifically, for a weight recovery attack, we set $\mu=0$ and $\sigma^{2}=0.5$ in Algorithm 1. In Scenario 1, $\alpha_{s}$ values were drawn from a normal distribution with a mean of 00 and a variance of $0.09$. For Scenarios 2 and 3, $\alpha_{s}$s values were generated from a normal distribution with a mean of 00 and a variance of $1$. Weight recovery in Scenarios 1 and 2 follows Algorithm 1 while Scenario 3 adheres to the algorithm with one modification: line 1 is replaced by
$\mathcal{H}[j-1]=\mathcal{L}(\hat{\boldsymbol{v}}[j-1]),$
in accordance with Equation 1 using the pre-generated leakage coefficients and noise.
For the recovery of weights, a total of $250$ attacks were conducted, with randomly generated weight values for each attack. Similarly, $250$ attacks were simulated for the recovery of biases, with randomly generated bias values for each attack.
@@ 4.2 Application to Full Networks
As can be seen from Table 2, we observe that the recovered model from Scenario 1 and 3 performs almost similarly to the original GoogleNet model. This makes sense since the leakage model quite matches the leakage behavior. However, for Scenario 2, when the attacker is using the Hamming weight leakage model when targeting leakage with high variance for their bit coefficients, the performance degrades significantly. This follows with the idea of the scenario which shows the worst case for the attacker, with random coefficients and no knowledge of them [^2] However, such a scenario would be rare in the real world..
@@ Acknowledgement
OpenAI’s ChatGPT-4 was used to improve the clarity and readability of this manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the manuscript’s content. This project has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under the Programme SASPRO 2 COFUND Marie Sklodowska-Curie grant agreement No. 945478. This research is funded by the European Commission, under the Horizon Europe project aerOS, grant number 101069732.
This research is supported by the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity Research & Development Programme (Development of Secured Components & Systems in Emerging Technologies through Hardware & Software Evaluation ¡ NRF-NCR25-DeSNTU-0001 ¿). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the view of National Research Foundation, Singapore and Cyber Security Agency of Singapore.